40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Feb 26 22:05:06 2020, gautam, Update, IOO, IMC checks 
    Reply  Wed Feb 26 23:50:51 2020, gautam, Update, IOO, IMC checks sliderCal.pdfAO_inputTFs.pdf
Message ID: 15227     Entry time: Wed Feb 26 22:05:06 2020     Reply to this: 15229
Author: gautam 
Type: Update 
Category: IOO 
Subject: IMC checks 

Today, I did the following tests (and so was touching electronics/cables at/around 1X2):

  • Measured the IMC OLTF.
  • Measured the TF from injection at IN2 to response at the IMC error point (T-eed the I out of the IMC demodulator as there is no longer a monitor point available).
  • Measured the IMC in loop error signal with 0.25 Hz resolution from DC-2kHz.
  • Confirmed that the IMC IN2 (a.k.a. AO path) gain slider performs as advertised. This is a useful test to run post Acromag switchover on Friday.

Results to follow.

After this work, I reverted the EPICS channels to the usual values. The IMC can be locked.

ELOG V3.1.3-